OpenStudy (anonymous):

what's the inverse of f(x)=2x-10 ?

OpenStudy (kropot72):

The inverse of a function is obtained by exchanging x and y in the equation of the function. f(x) = y y = 2x - 10.....................(1) x = 2y -10 (after exchanging x and y in (1) )................(2) 2y = x + 10 \[f ^{-1}(x)=\frac{x+10}{2}\]